MOLE CONCEPT Q: MnO2 on ignition converts into Mn3O4.A pyrolusite having 75% MnO2 ,20% inert impurities and rest water is ignited in air to a constant weight. what is the percentage of Mn in the ignited sample?  

Grade:

1 Answers

Sunil Kumar FP
9 years ago
apply POAC Mn atom
let there be 100g mixture
then wtg. of Mn= 75g now, inert impurities =20g and xg Mn3O4 is formed
MnO2 = Mn3O4
for Mn atom 75/87 =3x/229( M.W. of MnO2 = 87, M.W. of Mn3O4 =229)
then x=65.8
wt of mn in product = (75/87) * 55=47.41g
%tage of Mnin product = 47.41/(65.8+20) * 100=55.26%